Operation Summary for April 8, 2011 - April 11, 2011 |
The Integrated luminosity for the period from 4/4/11 to 4/11/11 was 62.89 pb-1 (inverse picobarns). NuMI took 1.16E18 of beam during this same time period. |
|
Friday April 8 The day shift began with Operations monitoring the Tevatron (TEV) store 8631 (155.42E30) and the Recycler stash (157.0E10), with the Antiproton Source (Pbar) stacking (23.4E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI experiment preparing to make target scans, with the MiniBooNE and Meson FTFBT-978 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, NuMI personnel conducted target scans for most of the shift and then switched to high intensity studies. MI personnel conducted G minus 2 (G -2) studies for a possible future experiment. During the evening shift, Operations terminated store 8631 at 5:34 PM. Operations disabled beam to T-978 at 6 PM, put the TEV into shot setup, and established store 8632 at 7:02 PM with an average initial luminosity (AIL) of 311.2E30. EE Support personnel replaced a blown fuse in the PLC for the MI-21 amplifiers. Operations monitored the NuMI power supplies, which were left pulsing after their studies had finished.Saturday April 9 The midnight shift began with Operations monitoring the store 8632 (141.08E30) and the stash (168.6E10), with Pbar stacking (11.0E10), and with MiniBooNE taking beam. During the midnight shift, Operations established beam to T-978 at 4:20 AM. Operators prepared MI-65 at 5:20 AM for an access. During the day shift, Operations terminated store 8632 at 10:47 AM, put the TEV into shot setup, and established store 8633 at 11:57 AM with an AIL of 333.1E30. Operations reported that the Pbar dump LCW system developed a leak that required filling. NuMI started taking beam at noon. Booster and MI personnel tuned their machines during most of this shift. During the evening shift, Operations disabled beam to T-978 at 4:35 PM. Sunday April 10 The midnight shift began with Operations monitoring store 8633 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations reported several hours of downtime when a Linac Klystron RF station (KRF1) had a charging supply failure that required repair; operators called in Linac technicians. Operators filled the Pbar dump LCW system several times this shift. Operations established beam to T-978 at 6:45 AM (after the repair of KRF1). During the day shift, Operations terminated store 8633 at 8:27 AM, put the TEV into shot setup, and established store 8634 at 9:47 AM with an AIL of 303.48E30. Operators tuned on Linac and Booster. During the evening shift, Operations disabled beam to T-978 at 6 PM. Monday April 11 The midnight shift began with the TEV in shot setup, with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 8634 at 2:10 AM, put the TEV into shot setup, and established store 8635 at 4:19 AM with an AIL of 310.7E30. Operations established beam to T-978 at 4:59 AM. Operations reported that a Linac RF station (LRF2) went down pretty hard at 6:53 AM; Operators reported hearing arcing from inside the anode cabinet. Operators called in the Linac On-Call RF tech to look into the problem. Operations reported minor Booster RF trips during this shift. The Plans for Monday The plans for today are to stack, stash, and store. The Run Coordinator will allow MI personnel to check coalescing and their MI-8 collimator positions. She will also allow Booster and Linac personnel to conduct studies. |
|
More Information Comments and Suggestions |
| last modified 04/11/2011 email Fermilab |
| Security, Privacy, Legal | |